OpenStudy (anonymous):

If A*B=A+2B, What is the value of 7*3?

OpenStudy (anonymous):

7*3=7+(2*3)==13

OpenStudy (skullpatrol):

$$\Huge 7*3=21$$

OpenStudy (skullpatrol):

With or without your "If" statement.

OpenStudy (alekos):

i think he means that the * is an operator not a multiplication